OpenStudy (dakid88):

An education budget went from $800,000 to $900,000. What is the approximate percent increase? The approximate percent increase is 13%. The approximate percent increase is 89%. The approximate percent increase is 98%. The approximate percent increase is 113%.

OpenStudy (dakid88):

@maddie82502

OpenStudy (dakid88):

@Lord2clash

OpenStudy (dakid88):

@ganeshie8

OpenStudy (dakid88):

anyone

OpenStudy (lord2clash):

well there was another question exactly like this one year ago so im gonna leave a link and you can look over that. Link : http://inquus.org/updates/53da4fd9e4b09f69dcc1e906

OpenStudy (lord2clash):

(900-800)/800*100 = 12.5 % or (900,000 - 800,000) / (800,000) 100,000/800,000 = 0.125 0.125 * 100 = 12.5%

OpenStudy (dakid88):

i know but 12.5% is not one of the answers @Lord2clash

OpenStudy (dakid88):

wht do u think the answer is @boolean

jigglypuff314 (jigglypuff314):

the question is asking for "approximate" so what is the answer closested to the one @Lord2clash got you? :)

OpenStudy (anonymous):

Alright, our starting value is: 800,000, the final value is 900,000 Percentage increase can be found by taking the difference between 900,000 and 800,000 So... 900,000 - 800,0000 = 100,000 We know now the increase was $100,000 Divide this value by the original/starting value ($800,000) and then * 100. \[(100,000 \div 800,000) \times 100 = 12.5\]

OpenStudy (dakid88):

k thxs

OpenStudy (dakid88):

can u give a medal too me too

jigglypuff314 (jigglypuff314):

for doing no work yourself? pluh-leaz you gotta earn it

OpenStudy (dakid88):

plz by

OpenStudy (dakid88):

can yahll help me with three more